A proton moves rectilinearly. Its movement (its position) is governed by the following equation: x (t) = 7.8 + 9.2 t - 2.1 t^3, where t is expressed in s.

When does this proton reach its maximum position in x?
Answer #1

x(t) = 7.8 + 9.2t - 2.1t^3

differentiate with respect to t,

d(x(t))/dt = d/dt(7.8 + 9.2t - 2.1t^3)

v(t) = 9.2 - 6.3t^2

put v(t) = 0

9.2 - 6.3t^2 = 0

t^2 = 9.2/6.3

t^2 = 1.460

t = 1.21 s

differentiate v(t) with respect to t,

d/dt(v(t)) = d/dt(9.2 - 6.3t^2)

= -12.6t

At t = 1.21 s,

d/dt(v(t)) is negative,

So, t = 1.21 s gives maximum value of x(t)

Answer: 1.21 s